# On the Fundamental Limitations of Dual Static CVaR Decompositions in Markov Decision Processes

### v1

@@ 5 From Impossible Evaluation to Impossible Uniform Optimality
Leveraging the risk-assignment constraints perspective developed in the previous section, we now show that there exists an MDP where it is impossible for a single risk-dependent policy $\tilde{\pi}$ to be uniformly optimal, that is, being optimal for all initial risk levels $\alpha\in[0,1]$ simultaneously. This suggests that the limitations of the dual CVaR decomposition are fundamental to the current problem formulation and cannot be solved by simple algorithmic improvements. We begin by formalizing the notion of a uniformly optimal policy.
Definition 10 (Uniformly optimal policy)
A risk-dependent policy $\tilde{\pi}$ is uniformly optimal if its corresponding history-dependent policy $\tilde{\pi}_{h}^{\alpha}$ is optimal for all initial risk levels $\alpha\in[0,1]$. That is, for all $\alpha\in[0,1]$, we have
@@
For a policy $\tilde{\pi}$ to achieve uniform optimality, selected actions must align with those of the true optimal history-dependent policy $\pi_{h,\alpha}^{\star}$ for every value of $\alpha$ and all possible histories. By Corollary 9, each optimal policy $\pi_{h,\alpha}^{\star}$ induces a unique risk-dependent $\tilde{\pi}_{\alpha}$ and a corresponding risk-level assignment $\mathcal{Y}_{\alpha}$. To be uniformly optimal, a policy $\tilde{\pi}$ therefore has to ensure it simultaneously follows every $\tilde{\pi}_{\alpha}$ alongside its risk-level assignment $\mathcal{Y}_{\alpha}$. These requirements can be grouped in a set of constraints, which we refer to as the optimal-action-selection constraints.
@@
[Uniform optimality constraints]propositionnpropositionConditionUniformOptimality A risk-dependent policy $\tilde{\pi}$ is uniformly optimal if and only if it simultaneously satisfies all the optimal-action-selection constraints
@@
defined for all initial risk-levels $\alpha\in[0,1]$, optimal policies $\pi_{h,\alpha}^{\star}$, histories $H\in\mathcal{H}_{T}$ with $P^{\pi_{h,\alpha}^{\star}}(H)>0$, and time steps $t=0,\dots,T-1$.
@@
Theorem 11
@@
There exists an MDP for which no single risk-dependent policy $\tilde{\pi}:\mathcal{S}\times[0,1]\to\mathcal{A}$ is uniformly optimal.
@@
Proof  We prove the result by providing an example MDP for which uniform optimality is impossible. We once again use the MDP from Hau et al. (2023), displayed in Figure 1, with horizon $T=2$. This MDP contains only three different history-dependent policies, which are all fully characterized by the selected action in state $s_{1}$. We can therefore easily compute the static CVaR of history-dependent policies to see which one is optimal at different initial risk-level $\alpha$. To simplify notation, let us denote the three history-dependent policies $\pi_{h}^{(i)}$ to indicate which action $a_{i}$ they select in $s_{1}$. Figure 3 shows the CVaRα of each history-dependent policy based on the initial risk-level $\alpha$. We can therefore deduce the optimal policy:
$\pi_{h,\alpha}^{\star}\coloneq\begin{cases}\pi_{h}^{(2)}&\text{if}\quad\alpha\in[0,0.375)\\ \pi_{h}^{(3)}&\text{if}\quad\alpha\in[0.375,0.6875)\\ \pi_{h}^{(1)}&\text{if}\quad\alpha\in[0.6875,1].\end{cases}$
@@
Figure 3: Evolution of the static CVaR evaluation of all possible policies on the MDP presented in Figure 1 at different initial risk levels $\alpha$. Shaded regions represent the optimal policy $\pi_{h}^{\star}$ at a given initial risk-level $\alpha$.
@@
Corollary 9 applies to every policy $\pi_{h}^{(i)}$ because they are risk-independent, hence computing their respective value functions (Eq. 6) grants us state-level perturbations $\bm{\tilde{\xi}}$ from which we can extract a consistent risk assignment. Combining these risk assignments with the $\alpha$ values where each $\pi_{h}^{(i)}$ is optimal, we can extract the optimal-action-selection constraints. To streamline our argument, let us consider the state-level risk envelope constraints imposed at state $s_{1}$ for cases $\alpha=0.25$, $\alpha=0.5$, and $\alpha=0.75$ in particular. For these, the optimal-action-selection-constraints in $s_{1}$ yield:
$\pi_{h,0.25}^{\star}=\pi_{2}$$\implies$ $\mathcal{Y}_{0.25}(s_{1})=0.5$$\implies$ $\tilde{\pi}^{\star}(s_{1},0.5)=a_{2}$
$\pi_{h,0.5}^{\star}=\pi_{3}$$\implies$ $\mathcal{Y}_{0.5}(s_{1})=0.5$$\implies$ $\tilde{\pi}^{\star}(s_{1},0.5)=a_{3}$
$\pi_{h,0.75}^{\star}=\pi_{1}$$\implies$ $\mathcal{Y}_{0.75}(s_{1})=0.5$$\implies$ $\tilde{\pi}^{\star}(s_{1},0.5)=a_{1}$
@@
That is, in order to be simultaneously optimal for initial risk levels $\alpha=0.25$, $\alpha=0.5$, and $\alpha=0.75$, a risk-dependent policy $\tilde{\pi}$ is required to take all actions $a_{1}$, $a_{2}$, and $a_{3}$ in state $S_{t}=s_{1}$ when the risk-level is $Y_{t}=0.5$, proving the impossibility of uniform optimality.
@@
Figure 4: Relation between initial risk-level $\alpha$ and the corresponding risk-level $\mathcal{Y}(s_{1})$ for all three possible policies on the MDP presented in Figure 1. Shaded regions represent the optimal policies, with the resulting constraint on $\tilde{\pi}$ displayed explicitly.
@@
Note that the range of values preventing the existence of a uniformly optimal policy in the MDP presented in Figure 1 extends beyond the $\alpha$ values presented in the proof of Theorem 11. In Figure 4, we visualize the optimal-action-selection constraints for all $\alpha\in[0,1]$. The figure highlights the presence of a wide range of $\alpha$ where the optimal-action-selection constraints overlap, precluding the presence of a single optimal risk-dependent policy $\pi$ simultaneously optimal for these initial risk-levels $\alpha$.
@@ 6 Conclusion
In this work, we diagnosed the root cause of failures in the static dual CVaR dynamic program decomposition. We started by framing the problem of evaluating a policy’s static CVaR as two distinct but related optimization tasks: one over history-level perturbations for the true static CVaR and another over state-level perturbations for the dynamic program. This perspective revealed that a CVaR evaluation gap arises precisely when a set of risk-assignment consistency constraints have an empty intersection. Building on this risk-assignment constraint perspective, we proved that the dual decomposition itself is fundamentally limited by identifying an MDP where no single risk-dependent policy can be uniformly optimal for all initial risk levels, as the action requirements for optimality at different risk levels become contradictory.
@@
Our findings show that seeking a single, uniformly optimal policy with the current dual decomposition approach is flawed. Future work should therefore pivot towards developing algorithms that find the optimal policy for a specific initial risk level, similar to methods used for primal-based decompositions (Bäuerle and Ott, 2011). Another alternative research direction is to identify the conditions on the MDP or the policy class under which the risk-assignment consistency constraints are always satisfiable, thereby guaranteeing the absence of an evaluation gap. Ultimately, our analysis provides a foundation for further characterizing, and potentially overcoming, the challenges of risk-averse reinforcement learning.
Acknowledgements
This work was supported by NSERC ES D scholarship 569096-22, FRQNT scholarship 318186, and a Canada-CIFAR AI Chair.
Appendix A Omitted proofs
@@
We now restate the results from the main body and present their omitted proofs.
A.1 Results of Section 3
@@
Proof  We prove the result by contradiction. Let $\bm{\tilde{\xi}}^{\star}\coloneq\mathop{\mathrm{arg\,min}}\limits_{\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}(\tilde{\pi})}\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s,y)$. Suppose there exists $\bm{\tilde{\xi}}^{\prime}$ and $(s,y)$ such that $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{T}(s,y)<\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T}(s,y)$. Then, by the definition of $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T}(s,y)$ (Eq. 8), we must have
$\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a)\bm{\tilde{\xi}}^{\prime}(s^{\prime}|s,y,a)\left(\mathcal{R}(s,a,s^{\prime})+\gamma\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{T-1}(s^{\prime},y^{\prime})\right)$
$<$ $\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a))\bm{\tilde{\xi}}^{\star}(s^{\prime}|s,y,a)\left(\mathcal{R}(s,a,s^{\prime})+\gamma\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T-1}(s^{\prime},y^{\prime})\right)$
$=$ $\min_{\tilde{\xi}\in\Xi_{y}(P(\cdot|s,a)}\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a)\tilde{\xi}(s^{\prime}|s,y,a)\left(\mathcal{R}(s,a,s^{\prime})+\gamma\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T-1}(s^{\prime},y^{\prime})\right),$
where we applied the definition of the dynamic program at step $T$ (Eq. 6) in the last equation. However, since $\min$ is monotonous with $P$ and $\mathcal{R}$ remaining constant, the above relation implies that the inequality also held at the previous time step $T-1$ for at least one $(s^{\prime},y^{\prime})$. That is, there exists $(s^{\prime},y^{\prime})$ such that $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{T-1}(s^{\prime},y^{\prime})<\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T-1}(s^{\prime},y^{\prime})$. Repeating this argument $T$ times, this implies that there exists $(s,y)$ such that $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{0}(s,y)<\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{0}(s,y)$, which is impossible because they all are initialized to 0.

### v2

@@ 5 From Impossible Evaluation to Impossible Uniform Optimality
One of the most significant consequence of using the dual-based CVaR decomposition (Pflug and Pichler, 2016) and finding a single optimal risk-dependent policy $\tilde{\pi}$ is that it could simultaneously retrieve the optimal history-dependent policy for any initial risk-level (Chow et al., 2015), a property we will refer to as uniformly optimality. This property represented a massive advantage over the primal-based decomposition (Bäuerle and Ott, 2011), which is constrained to find the optimal policy for a single initial risk-level. Leveraging the risk-assignment constraints perspective developed in the previous section, we now show that there exists an MDP where it is impossible for a single risk-dependent policy $\tilde{\pi}$ to be uniformly optimal, proving this inherent property to be impossible in general. We begin by formalizing the notion of uniform optimality.
Definition 5 (Uniformly optimal policy).
A risk-dependent policy $\tilde{\pi}$ is uniformly optimal if its corresponding history-dependent policy $\tilde{\pi}_{h}^{\alpha}$ is optimal for all initial risk levels $\alpha\in(0,1]$. That is, for all $\alpha\in(0,1]$, we have
@@
For a policy $\tilde{\pi}$ to achieve uniform optimality, its selected actions must align with those of the optimal history-dependent policy $\pi_{h,\alpha}^{\star}$ for every value of $\alpha$ and all possible histories. By Corollary 4, each optimal policy $\pi_{h,\alpha}^{\star}$ induces a unique risk-dependent policy $\tilde{\pi}_{\alpha}$ and risk level assignment $\mathcal{Y}_{\alpha}$. To be uniformly optimal, a policy $\tilde{\pi}$ therefore has to ensure it simultaneously follows every $\tilde{\pi}_{\alpha}$ alongside its risk level assignment $\mathcal{Y}_{\alpha}$. These requirements can be grouped in a set of optimal-action-selection constraints.
@@
[Uniform optimality constraints]propositionpropositionConditionUniformOptimality A risk-dependent policy $\tilde{\pi}$ is uniformly optimal if and only if it simultaneously satisfies all the optimal-action-selection constraints
@@
defined for all initial risk levels $\alpha\in(0,1]$, optimal policies $\pi_{h,\alpha}^{\star}$, histories $H\in\mathcal{H}_{T}$ with $P^{\pi_{h,\alpha}^{\star}}(H)>0$, and time steps $t=0,\dots,T-1$.
@@
Theorem 2 (Impossible uniform optimality counterexample).
@@
There exists an MDP satisfying Assumptions 1 for which no single risk-dependent policy $\tilde{\pi}:\mathcal{S}\times(0,1]\to\mathcal{A}$ is uniformly optimal.
Proof.
@@
We prove the result by providing an example MDP for which uniform optimality is impossible. We once again use the MDP from Hau et al. (2023), displayed in Figure 1, with horizon $T=2$. This MDP contains only three different history-dependent policies, which are all fully characterized by the selected action in state $s_{1}$. We can therefore easily compute the static CVaR of history-dependent policies to see which one is optimal at different initial risk level $\alpha$. To simplify notation, let us denote the three available policies $\pi_{h}^{(i)}$ to indicate which action $a_{i}$ they select in $s_{1}$.
Figure 3 shows the CVaRα of each policy based on the initial risk level $\alpha$. Crucially, Figure 3 shows that the optimal action choice is strictly dependent on the initial risk level. This observation serves as the foundational mechanism for our counterexample, illustrating how action requirements for optimality at different risk levels eventually become contradictory. We can deduce the optimal policy:
$\pi_{h,\alpha}^{\star}\coloneq\begin{cases}\pi_{h}^{(2)}&\text{if}\quad\alpha\in[0,0.375)\\ \pi_{h}^{(3)}&\text{if}\quad\alpha\in[0.375,0.6875)\\ \pi_{h}^{(1)}&\text{if}\quad\alpha\in[0.6875,1],\end{cases}$
meaning that the optimal risk-seeking ($\alpha$ close to 1) choice is to select action $a_{1}$, while the optimal risk-averse ($\alpha$ close to 0) choice is instead to pick $a_{2}$, with $a_{3}$ being the optimal choice at a moderate risk-level.
@@
Figure 3: Evolution of the static CVaR evaluation of all policies on the MDP presented in Figure 1 at different initial risk levels $\alpha$. Shaded regions represent the optimal policy $\pi_{h}^{\star}$ at a given initial risk level $\alpha$.
@@
Because all three policies $\pi_{h}^{(i)}$ are Markovian, Corollary 4 applies and we know that computing their respective value functions (Eq. 7) grants us state-level perturbations $\bm{\tilde{\xi}}$ from which we can extract a consistent risk assignment. Combining these risk assignments with the $\alpha$ values where each $\pi_{h}^{(i)}$ is optimal, we can extract the optimal-action-selection constraints. To streamline our argument, let us consider the state-level risk envelope constraints imposed at state $s_{1}$ for cases $\alpha=0.25$, $\alpha=0.5$, and $\alpha=0.75$ in particular. For these, the optimal-action-selection-constraints (Proposition 5) in $s_{1}$ yield:
$\pi_{h,0.25}^{\star}=\pi_{h}^{(2)}$ $\implies$ $\mathcal{Y}_{0.25}(s_{1})=0.5$ $\implies$ $\tilde{\pi}^{\star}(s_{1},0.5)=a_{2}$
$\pi_{h,0.5}^{\star}=\pi_{h}^{(3)}$ $\implies$ $\mathcal{Y}_{0.5}(s_{1})=0.5$ $\implies$ $\tilde{\pi}^{\star}(s_{1},0.5)=a_{3}$
$\pi_{h,0.75}^{\star}=\pi_{h}^{(1)}$ $\implies$ $\mathcal{Y}_{0.75}(s_{1})=0.5$ $\implies$ $\tilde{\pi}^{\star}(s_{1},0.5)=a_{1}$
@@
That is, in order to be simultaneously optimal for initial risk levels $\alpha=0.25$, $\alpha=0.5$, and $\alpha=0.75$, a risk-dependent policy $\tilde{\pi}$ would be required to take all actions $a_{1}$, $a_{2}$, and $a_{3}$ in state $S_{t}=s_{1}$ when the risk level is $Y_{t}=0.5$, proving the impossibility of uniform optimality. ∎
@@
Figure 4: Relation between initial risk level $\alpha$ and the corresponding risk level $\mathcal{Y}_{\alpha}(s_{1})$ for all three possible policies on the MDP presented in Figure 1. Colored areas indicate the optimal policies, with the resulting constraint on $\tilde{\pi}$ displayed explicitly.
@@
Note that the range of values preventing the existence of a uniformly optimal policy in the MDP presented in Figure 1 extends beyond the $\alpha$ values presented in the proof of Theorem 2. To illustrate this, Figure 4 displays the optimal-action-selection constraints for all $\alpha\in(0,1]$. Colored regions indicate which policy is considered optimal for the given initial risk level $\alpha$, with a box highlighting the explicit constraint for every region. For every policy, the risk-assignment in $s_{1}$ ($\mathcal{Y}_{\alpha}(s_{1})$) it corresponds to for an initial risk-level $\alpha$ is shown in a separate color. The figure highlights the presence of a wide range of initial risk levels $\alpha$ where the optimal-action-selection constraints overlap, precluding the presence of a single optimal risk-dependent policy $\pi$ simultaneously optimal for these initial risk levels.
@@ 6 Conclusion
In this work, we diagnosed the root cause of failures in the static dual CVaR dynamic program decomposition. We started by framing the problem of evaluating a policy’s static CVaR as two distinct but related optimization tasks: one over history-level perturbations for the true static CVaR and another over state-level perturbations for the dynamic program. This perspective revealed that a CVaR evaluation gap arises precisely when a set of risk-assignment consistency constraints (Def. 4) have an empty intersection (Thm. 4). Our findings provide another independent proof that Markovian policies enjoy a null CVaR evaluation gap, which is unfortunately not the case for risk-dependent policies. Building on our risk-assignment constraints, we proved that the dual decomposition itself is fundamentally limited by identifying an MDP where no single risk-dependent policy can be uniformly optimal for all initial risk levels, as the action requirements for optimality at different risk levels become contradictory (Thm. 2).
Future Directions
@@
Our findings show that seeking a single, uniformly optimal policy with the current dual decomposition approach is flawed because it can involve incompatible constraint sets. Future work should therefore pivot towards developing algorithms that find the optimal policy for a specific initial risk level, similar to methods used for primal-based decompositions (Bäuerle and Ott, 2011). Even when limiting our attention to a single initial risk level, many interesting questions surrounding proper evaluation remain.
  1. 1.
    What is the extent of risk-dependent policies that have a CVaR evaluation gap at any given initial risk-level? On one hand, we presented an instance where a risk-dependent policy presented a CVaR evaluation gap (Sec. 4). On the other hand, one can also easily find examples that do not suffer from a CVaR evaluation gap (App. C). This leaves open the characterization of the occurrence of evaluation gaps. For instance, can we always find a corresponding risk-dependent policy without evaluation gap for any history-dependent policy and initial risk-level? Is this characterization dependent on intrinsic properties of the MDP or universally applicable?
  2. 2.
    Can we determine if a risk-dependent policy will have a CVaR evaluation gap without access to its corresponding history-level perturbations? The risk assignment consistency constraints (Def. 4) rely on knowledge of the corresponding history perturbations. This is problematic in practice as it requires the standalone static evaluation of a risk-dependent policy’s history-dependent counterpart to evaluate whether the risk-level decomposition is accurate. One possible path forward we envision is to leverage the set of all risk-dependent policies that share the same history-dependent counterpart for a given initial risk-level. For instance, should we know that at least one of these policies does not have an evaluation gap, then as a consequence of the static CVaR upper-bound (Cor. 1), we would know that the risk-dependent policies with lowest reported DP value would be without CVaR evaluation gap.
Ultimately, despite the identified limitations regarding uniform optimality, the static CVaR dual decomposition retains significant practical appeal. In contrast to primal-based methods (Bäuerle and Ott, 2011), which necessitate augmenting states with unbounded cumulative returns, the dual formulation operates within the bounded $(0,1]$ risk interval. This boundedness allows for a standardized, domain-independent discretization strategy that is more tractable in practice. Our analysis provides a foundation for further characterizing, and potentially overcoming, the challenges of risk-averse reinforcement learning. By formally characterizing the evaluation gap, our work identifies the precise constraints that any valid solution must satisfy. This analysis provides the necessary theoretical pathway for future research to derive exact, target-specific dual algorithms that reconcile these computational benefits with rigorous optimality guarantees.
Appendix A Notation
Symbol Description
$\mathcal{S}$ Finite state space
$\mathcal{A}$ Finite action space
$\Delta(\cdot)$ Probability simplex
$P$ Transition function $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$
$\mathcal{R}$ Reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,R_{\text{max}}]$
$\gamma$ Discount factor
$s_{0}$ Initial state
$T$ Trajectory horizon
$H$ History $(S_{0},A_{0},S_{1},\dots)$
$\mathcal{H}_{t}$ Set of histories of length $t$
$H_{0:k}$ $k$-length subsequence of history $H$
$\pi_{h}$ History-dependent policy $\pi_{h}:\mathcal{H}\mapsto\mathcal{A}$
$\pi$ Markovian policy $\pi:\mathcal{S}\mapsto\mathcal{A}$
$P^{\pi_{h}}(H)$ Probability of history $H$ when following policy $\pi_{h}$
$\mathcal{R}^{H}_{0:T}$ Return of history $H$
$Z^{\pi_{h}}$ Random return when sampling histories by following policy $\pi_{h}$
$\Xi_{\alpha}(\mathbb{P})$ CVaRα risk envelope around probability distribution $\mathbb{P}$
$\xi$ History-level perturbations $\xi:\mathcal{H}\mapsto[0,\frac{1}{\alpha}]$, $\xi\in\Xi_{\alpha}(P^{\pi})$
$Y$ Running risk level $Y\in(0,1]$
$\tilde{\mathcal{S}}$ Risk-augmented state space $\tilde{\mathcal{S}}\coloneq\mathcal{S}\times(0,1]$
$\tilde{\pi}$ Risk-dependent policy $\tilde{\pi}:\tilde{\mathcal{S}}\mapsto\mathcal{A}$
$\bm{\tilde{\xi}}$ State-level perturbations $\bm{\tilde{\xi}}(s,a,y)\in\Xi_{y}(P(\cdot|s,a))$
$\bm{\tilde{\Xi}}$ All valid state-level perturbations $\bm{\tilde{\Xi}}\coloneq\left\{\bm{\tilde{\xi}}|\newline \,\bm{\tilde{\xi}}(\cdot|s,y,a)\in\Xi_{y}\left(P(\cdot|s,a)\right)\forall(s,y,a)\in\mathcal{S}\times(0,1]\times\mathcal{A}\right\}$
$\tilde{\pi}_{h}^{\alpha}$ History-dependent correspondance of risk-dependent policy $\tilde{\pi}$ at initial risk level $\alpha\in(0,1]$
$\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}$ Policy-perturbations value function $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}:\mathcal{S}\times\mathcal{(}0,1]\mapsto\mathbb{R}$
$\zeta_{\alpha}^{\bm{\tilde{\xi}}}$ Perturbation mapping $\zeta_{\alpha}^{\bm{\tilde{\xi}}}\in\Xi_{\alpha}(P^{\pi_{h}})$
$\mathcal{Y}$ risk level assignment $\mathcal{Y}:\mathcal{H}\mapsto(0,1]$
Table 1: List of notations
Appendix B Omitted proofs
@@
We now restate the results from the main body and present their omitted proofs. For every result, we also include a proof outline to provide a clearer picture of the proof’s content.
B.1 Results of Section 3
@@
Proof outline.
The result is proven by contradiction. We assume that there exist state-level perturbations yielding a value strictly lower than the minimum defined by the dynamic program at the final time step at any state-risk level pair $(s,y)$. We then demonstrate that due to the recursive nature of the value function (Def. 2), this strict inequality must propagate backwards through time, ultimately implying that the inequality must hold at time $t=0$, contradicting the uniform initialization of the value function to zero.
Proof.
We prove the result by contradiction. First observe that by definition of $\bm{\upsilon}^{\tilde{\pi}}$, there must exist $\bm{\xi}\in\bm{\tilde{\Xi}}$ such that $\bm{\upsilon}^{\tilde{\pi}}\coloneq\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}$. Let $\bm{\tilde{\xi}}^{\star}\coloneq\mathop{\mathrm{arg\,min}}\limits_{\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}}\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s,y)$ and suppose there exists $\bm{\tilde{\xi}}^{\prime}\in\bm{\tilde{\Xi}}$ and $(s,y)$ such that $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}(s,y)<\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}(s,y)$. Then, by the definition of $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}(s,y)$ (Eq. 9), we must have
$\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{T}(s,y)<\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T}(s,y).$
Let $a=\tilde{\pi}(s,y)$. Expanding the LHS using the definition of the policy-perturbations value function (Eq. 9) and the RHS using the dynamic program definition (Eq. 7), we get
$\,\sum_{s^{\prime}}P(s^{\prime}|s,a)\bm{\tilde{\xi}}^{\prime}(s^{\prime}|s,y,a)\left[\mathcal{R}(s,a,s^{\prime})+\gamma\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{T-1}(s^{\prime},y^{\prime})\right]$
$<$ $\,\min_{\tilde{\xi}\in\Xi_{y}\left(P(\cdot|s,a)\right)}\sum_{s^{\prime}}P(s^{\prime}|s,a)\tilde{\xi}(s^{\prime})\left[\mathcal{R}(s,a,s^{\prime})+\gamma\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T-1}(s^{\prime},y^{\prime})\right].$
Since $\bm{\tilde{\xi}}^{\prime}(\cdot|s,y,a)\in\Xi_{y}\left(P(\cdot|s,a)\right)$, the above minimum is upper bounded by substituting $\bm{\tilde{\xi}}^{\prime}$ directly:
$\,\min_{\tilde{\xi}\in\Xi_{y}\left(P(\cdot|s,a)\right)}\sum_{s^{\prime}}P(s^{\prime}|s,a)\tilde{\xi}(s^{\prime})\left[\mathcal{R}(s,a,s^{\prime})+\gamma\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T-1}(s^{\prime},y^{\prime})\right]$
$\leq$ $\,\sum_{s^{\prime}}P(s^{\prime}|s,a)\bm{\tilde{\xi}}^{\prime}(s^{\prime}|s,y,a)\left[\mathcal{R}(s,a,s^{\prime})+\gamma\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T-1}(s^{\prime},y^{\prime})\right].$
Combining these inequalities and canceling the common terms, we obtain
$\sum_{s^{\prime}}P(s^{\prime}|s,a)\bm{\tilde{\xi}}^{\prime}(s^{\prime}|s,y,a)\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{T-1}(s^{\prime},y^{\prime})<\sum_{s^{\prime}}P(s^{\prime}|s,a)\bm{\tilde{\xi}}^{\prime}(s^{\prime}|s,y,a)\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{T-1}(s^{\prime},y^{\prime}).$
For the above inequality to hold, there must exist at least one $(s^{\prime},y^{\prime})$ such that $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{T-1}(s^{\prime},y^{\prime})<\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}^{\star}}}_{T-1}(s^{\prime},y^{\prime})$.
Repeating this argument $T$ times implies the existence of $(s,y)$ such that $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}_{0}(s,y)<\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}_{0}(s,y)$, which is impossible because they all are initialized to 0. ∎